In an economy with no exports and​ imports, autonomous consumption is ​$3 ​trillion, the marginal propensity to consume
is 0.6​, investment is ​$5 ​trillion, and government expenditure on goods and services is ​$6 trillion. Taxes are ​$5 trillion and
do not vary with real GDP.
If real GDP is ​$34.4 ​trillion, calculate disposable​ income, consumption​ expenditure, and aggregate planned expenditure. What is equilibrium​ expenditure?
1
Expert's answer
2018-05-13T12:45:07-0400
Ca = $3 trillion, c = 0.6, I = $5 trillion, G = $6 trillion, T = $5 trillion, real GDP = $34.4 trillion. Disposable income Yd = Y - T = 34.4 - 5 = $29.4 trillion.
Consumption expenditure C = Ca + c*Yd = 3 + 0.6*29.4 = $20.64 trillion.
Aggregate planned expenditure equals real GDP, so it is $34.4 trillion.
Equilibrium expenditure AE = C + I + G = 20.64 + 5 + 6 = $31.64 trillion.